nomurphy496:

Rob Retiree gets a Social Security benefit of $800 per month. His wife receives a benefit equal to 50 percent of Rob's benefit. If they take an additional $300 per month from savings, what will their monthly income be?

jhonyy9:

his wife receives 50% of Rob's benefit - so how many will be this ? 50/100 = ? like a first step

mrinmo:

Rob gets $800 per month for social security benefits. His wife gets per month for social security benefits: \[\$ (50 \% \times 800) = \$ ( \frac{50}{100} \times 800) = \$ 400\] From their saving, they get $300 per month. Hence: Their total monthly family income is: \[ \$ ( 800 + 400 + 300 ) = \$ 1500 \]